OpenStudy (anonymous):

Find the surface area of the sphere with the given dimension. Leave your answer in terms of pi. 1.) Radius of 60m. 2.) Diameter of 14 cm.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (jdoe0001):

http://0.tqn.com/d/math/1/0/I/F/spherer.gif

OpenStudy (anonymous):

still need help?

OpenStudy (hunus):

If the formula for the surface area of a sphere is S.A. = 4*pi*r^2 what would the surface area be?

OpenStudy (anonymous):

hold on. doing other work. brb

OpenStudy (anonymous):

what is r?

OpenStudy (hunus):

r is the radius of the sphere

OpenStudy (anonymous):

I got 48,960...:/

OpenStudy (anonymous):

oh wait. 45,216.

OpenStudy (hunus):

That's correct, don't forget to include your units

OpenStudy (anonymous):

the answers are in pi

OpenStudy (anonymous):

and none are 45,000

OpenStudy (hunus):

Well then it will be (4*r^2)*pi

OpenStudy (hunus):

What's 4*60^2 ?

OpenStudy (anonymous):

14,400. :) Thanks.

OpenStudy (hunus):

Yup :)

OpenStudy (anonymous):

is it the same equation for the second one?

OpenStudy (hunus):

In the second one you're given the diameter of the sphere which is double the radius

OpenStudy (anonymous):

so, what would the equation be?

OpenStudy (hunus):

It would be the same equation just divide the diameter by two to get your radius

OpenStudy (anonymous):

so, instead of putting in 14, it would be 7?

OpenStudy (anonymous):

Can you help with 3 more please? :)

OpenStudy (anonymous):

Find the volume of the sphere shown. Give each answer rounded to the nearest cubic unit.

OpenStudy (hunus):

The equation for the volume of a sphere is \[\Large V = \frac{4\pi r^3}{3}\]

OpenStudy (anonymous):

Would the equation be the same for this one?

OpenStudy (hunus):

Yes, but you need to half the diameter again to get the radius

OpenStudy (anonymous):

okay, what about this one: "The volume of a sphere is 5,000pi m^3. What is the surface area of the sphere to the nearest square meter?"

OpenStudy (hunus):

Well if the surface area is 5000pi and we know that the equation for the volume of a sphere is 4/3 * pi * r^3 we need to set them equal to each other and solve for the radius (so we can use it to find the volume)

OpenStudy (anonymous):

how?

OpenStudy (hunus):

\[\Large 5000 \pi = 4\pi r^2\]

OpenStudy (anonymous):

314,000,000? :/

OpenStudy (hunus):

First you should cancel pi from both sides

OpenStudy (anonymous):

k

OpenStudy (anonymous):

then what?

OpenStudy (hunus):

Then we will have 4r^2 = 5000 What should you do next to have only r on one side

OpenStudy (anonymous):

get rid of the r by the 4?

OpenStudy (anonymous):

@Hunus

OpenStudy (hunus):

We need to divide the 4 to get rid of it

OpenStudy (anonymous):

so, it would be "2^2*5000"?

OpenStudy (hunus):

\[\frac{4r^2}{4}=\frac{5000}{4}\]

OpenStudy (anonymous):

what next?

OpenStudy (hunus):

If we have what r^2 equals what should we do to both sides to get r

OpenStudy (anonymous):

cross multiply?

OpenStudy (hunus):

What about the square root of r and the square root of 5000?

OpenStudy (anonymous):

ugh. I'm stuck.

OpenStudy (hunus):

If r^2 = 5000 then \[\sqrt{r^2} = r = \sqrt{5000}\]

OpenStudy (anonymous):

then what??

OpenStudy (anonymous):

@Hunus

OpenStudy (anonymous):

@Nurali please help

OpenStudy (anonymous):

never mind. I got it

OpenStudy (nurali):

ok